Suppose a monopoly can produce any level of output it wishes at a constant marginal and average cost of $5 per unit. Assume the monopoly sells its good in two separate markets. The demand curve of the first market is q1= 55-p1 and the demand curve of the second market is q2= 70-2p2
1
Expert's answer
2018-11-28T13:00:35-0500

P1=55-q1

TR1=55q1-q1*q1

MR1=55-2*q1

55-2q1=5

Q1=25

P1=55-25=30

P2=35-0.5q2

TR2=35q2-0.5q2*q2

MR2=35-q2

5=35-q2

Q2=30

P2=35-0.5*30=20